# Second- and third-order properties of multidimensional Langevin equations

### v13

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
where the time-derivative is interpreted in the Itô sense, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the covariation. In fact, the above follows simply from a rearrangement of terms:
@@
upon dividing by $\tau>0$ and taking $\tau\to 0^{+}$, where we used $\langle\mathbf{x}(\tau)\mathbf{x}(\tau)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Eq. (5) follows if the derivative can be taken inside the expectation (however, see Appendix B for a counterexample). This type of manipulation will become useful later.
@@
(9) $\mathbf{v}^{\mathsf{H}}\mathbf{D}\mathbf{v}=-\mathbf{v}^{\mathsf{H}}\left({\frac{\mathbf{A}+\mathbf{A}^{\mathsf{T}}}{2}}\right)\mathbf{v}=-(\Re\lambda)(\mathbf{v}^{\mathsf{H}}\mathbf{v})$
($\mathsf{H}$ denoting Hermitian conjugate and $\Re$ denoting the real part), therefore $-\Re\lambda$ is bounded by the smallest and largest eigenvalues of $\mathbf{D}$ [14].
@@ 2.2. Comparison of experimental data with theoretical predictions
Now, we turn to the question of comparing experimental data and theoretical predictions under a linear model. As stated in the introduction, we may not necessarily care about statistically resolvable deviations if they are quantitatively small. We thus desire a way to evaluate the size of deviations. We may consider comparing the deviation between experiment and theory of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ to $\mathbf{C}$. That is, if we consider a whitened dataset for which $\mathbf{C}=\mathds{1}$, the entries of the “deviation matrix” (i.e., experiment minus theory of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$) in these coordinates should be compared to unity. For a coordinate-independent description, we may look at the symmetric and antisymmetric parts of the deviation matrix and multiply these on the right by $\mathbf{C}^{-1}$, and inspect the eigenvalues, which are purely real and imaginary in the symmetric and antisymmetric cases, respectively.
We denote the deviation $\mathbf{M}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle_{\text{expt}}-\mathbf{C}_{\text{theo}}$, where “expt” and “theo” subscripts denote experimental and theoretical values, respectively. From now on we suppress the “theo” subscript and treat $\mathbf{C}$ as fixed. We may consider the eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$, which are real because we have a matrix (again, a second-rank mixed tensor) that is symmetric in coordinates in which $\mathbf{C}=\mathds{1}$. We may imagine an ensemble of stochastic systems, each having its own value of $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ which may differ from $\mathbf{C}$, and again denote the difference by $\mathbf{M}$. We assume that the ensemble mean of $\mathbf{M}$ is zero, and we now ask the question of what a suitable ensemble covariance of $\mathbf{M}$ would be that reflects significant deviations. We may then compare the elements of the experimental value of $\mathbf{M}$ to the square root of the ensemble variance. For biological applications, a factor of 0.3 or more may be considered significant.
@@
where $\sim$ denotes ensemble expectation. We may treat $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$, and multiply the l.h.s. (left-hand side) by the matrix inverse of the r.h.s. (right-hand side). We denote by $\bm{\Gamma}$ the resulting matrix:
@@
It so happens that $\bm{\Gamma}$ has at most one non-zero eigenvalue $\mu$, which is equal to its trace. (This is because in coordinates where $\mathbf{C}=\mathds{1}$, the matrix $\bm{\Gamma}$ takes the form $\mathbf{v}\mathbf{v}^{\mathsf{T}}$, and therefore has rank at most 1.) We have:
(12) $\operatorname{tr}(\bm{\Gamma})=M^{ij}M^{i^{\prime}j^{\prime}}C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}=\operatorname{tr}((\mathbf{M}\mathbf{C}^{-1})^{2}).$
Operating in “covariance-identity” coordinates, it is seen that $M^{ij}$ where $i\neq j$ is counted double as compared to when $i=j$. We give an example of this in two dimensions. Changing coordinates to $x^{\prime}:=(x+y)/\sqrt{2}$ and $y^{\prime}=(y-x)/\sqrt{2}$, we see that:
(13) $\langle{x^{\prime}}^{2}\rangle$ $=1+\langle xy\rangle,$
(14) $\langle{y^{\prime}}^{2}\rangle$ $=1-\langle xy\rangle,$
(15) $\langle x^{\prime}y^{\prime}\rangle$ $=0.$
This shows indeed that deviations in $\langle xy\rangle$ should be counted double compared to deviations in $\langle x^{2}\rangle$ or $\langle y^{2}\rangle$.
There is, however, a problem with Eq. (10), which is that it does not obey symmetry in the indices. We may denote by $v_{ij}$ a left eigenvector of $\bm{\Gamma}$ with eigenvalue $\mu$, which satisfies:
(16) $v_{ij}M^{ij}M^{i^{\prime}j^{\prime}}=\mu v_{ij}C^{ii^{\prime}}C^{jj^{\prime}}.$
@@
(17) $v_{ij}M^{ij}M^{i^{\prime}j^{\prime}}=\mu v_{ij}\frac{C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}}}{2}.$
Now if we neglect permutations in $(i,j)$ and $(i^{\prime},j^{\prime})$ [^5] I.e., only one of $(i,j)$ and $(j,i)$ is included in the list of “super-indices” (and similarly for $(i^{\prime},j^{\prime})$). This is necessary to invert $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$, since this quantity is invariant upon swapping $i\leftrightarrow j$ (or $i^{\prime}\leftrightarrow j^{\prime}$)., we can see that the matrix product of $M^{ij}M^{i^{\prime}j^{\prime}}$ with the matrix inverse of $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$ has left eigenvector with elements $v_{ij}+v_{ji}$ for $i\neq j$ and $v_{ij}$ for $i=j$ and eigenvalue $\mu$. Thus we may consider instead of Eq. (10) the relation:
@@
(19) $\operatorname{tr}(\bm{\Gamma})=\mu\sim\frac{d(d+1)}{2},$
where $d$ is the dimension. The r.h.s. is the number of degrees of freedom in $\mathbf{M}$, accounting for symmetry. If $\sqrt{\mu}$ is comparable to $\sqrt{d(d+1)/2}$, then it may be considered quantitatively significant.
Now we turn to covariance functions with a time-lag. We may again denote the deviation $\mathbf{M}:=\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{expt}}-\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{theo}}$. We can separate $\mathbf{M}$ into symmetric and antisymmetric components, denoted by $\mathbf{S}$ and $\mathbf{T}$ respectively:
@@
Straightforward calculation reveals that:
@@
The above equality relates its r.h.s. to the sums of the squared eigenvalues of $\mathbf{S}\mathbf{C}^{-1}$ and $\mathbf{T}\mathbf{C}^{-1}$. Since there is no symmetry requirement for $\mathbf{M}$, we may take the components of $\mathbf{M}$ as obeying Eq. (10), in which case the components of $\mathbf{S}$ and $\mathbf{T}$ follow:
@@
This implies:
@@
The r.h.s.’s correspond to the degrees of freedom in $\mathbf{S}$ and $\mathbf{T}$, respectively. If $\sqrt{M^{ij}M^{i^{\prime}j^{\prime}}C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}}$ is comparable to $d$, then $\mathbf{M}$ may be considered (collectively) quantitatively significant.
@@
2.3. Probability current density and angular momentum
The dynamics of the linear system is determined by the matrices $\mathbf{A}$ and $\mathbf{D}$. However, to alternatively characterize the system by quantities even and odd under time reversal, we may consider the angular momentum matrix:
@@ 2.3. Probability current density and angular momentum
Now we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. The angular momentum is $L:=\langle x\dot{y}-\dot{x}y\rangle$. It is related to the matrix $\bm{\Omega}$ by the relation
@@
where we used the Lyapunov relation. Upon taking determinants, we get
@@
In the three-dimensional case, $\bm{\Omega}$ has at most a single pair of non-zero purely imaginary eigenvalues, and the stochastic rotation frequency is half the magnitude of the angular momentum vector evaluated in “covariance-identity” coordinates.
The stochastic rotation frequency where $\mathbf{A}$ has two real eigenvalues has been computed in [15] and is easily derived from the above formula in coordinates where:
@@
At this point, we still do not have a good dimensionless measure of when broken detailed balance is quantitatively significant. To do this, we turn to a modified definition of gain matrix [13, 4]:
@@
where the last equality follows from the Lyapunov equation. The entropy production rate $\dot{S}$ is related to $\mathbf{H}$ as $\dot{S}=-\operatorname{tr}(\mathbf{A}\mathbf{H})$ [13]. By a similar argument as in [13], in coordinates in which $\mathbf{D}=\mathds{1}$, we see that $\mathbf{H}$ is antisymmetric and therefore, like $\bm{\Omega}$, has purely imaginary complex conjugate eigenvalues. As an alternative derivation, following [3], we may consider that $\mathbf{D}^{-1/2}\mathbf{H}\mathbf{D}^{1/2}$ is antisymmetric and therefore $\mathbf{H}$ is similar to an antisymmetric matrix. In two dimensions the (dimensionless) eigenvalues ±$\pm ih$ are given by a simple formula similar to that of $\omega_{\mathrm{stoch}}$:
@@
If $h$ is at least comparable to unity, then we consider that broken detailed balance is significant. (We may understand the factor of 2 in the definition of $\mathbf{H}$ by considering that the symmetric part of $\mathbf{A}\mathbf{C}$ is $-\mathbf{D}$ by the Lyapunov equation, so the antisymmetric part should be compared to $\mathbf{D}$.) We may illustrate this on the two examples above. For two real eigenvalues $-\lambda_{x}$ and $-\lambda_{y}$ of $\mathbf{A}$,
@@
For the case of complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ with $D^{xy}=0$,
@@
We see that $h\geq\omega/\lambda$. Again, there are two possibilities for significant broken detailed balance. One possibility is that $\omega/\lambda$ is significant compared to unity. The other possibility is $\sqrt{D^{xx}}\ll\sqrt{D^{yy}}$ (or vice versa), which allows for the possibility that $\omega\ll\lambda$. As before, this corresponds to the case where $\mathbf{A}$ almost has a generalized eigenvector.
For angular momenta of $d$-dimensional systems, we may follow the part of the previous discussion pertaining to the antisymmetric part, except with $\mathbf{C}$ replaced by $2\mathbf{D}$, i.e., for the “ensemble covariance”:
@@
Note that when summing the squared eigenvalues of $\mathbf{H}$, the eigenvalues occur in pairs. If the time-step is not small and angular momenta are measured using discrete time, then on the r.h.s. of the above, it may be suitable to use instead discrete-time estimators of the diffusion.

### v14

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
where the time-derivative is interpreted in the Itô sense, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the covariation defined by:
(6) $[x,y](t)=\lim_{\Delta\to 0}\sum_{n=0}^{N-1}(x(t_{n+1})-x(t_{n}))(y(t_{n+1})-y(t_{n}))$
where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=t$ and $\Delta:=\max\limits_{0\leq n\leq N-1}(t_{n+1}-t_{n})$. Its time derivative is given by:
(7) $\frac{\mathrm{d}[x,y](t)}{\mathrm{d}t}=\lim_{\tau\to 0^{+}}\frac{(x(t+\tau)-x(t))(y(t+\tau)-y(t))}{\tau},$
[…]
@@
upon dividing by $\tau>0$ and taking $\tau\to 0^{+}$, where we used $\langle\mathbf{x}(\tau)\mathbf{x}(\tau)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(0)^{\mathsf{T}}\rangle$ and took the derivative inside the expectation (however, see Appendix B for a counterexample). This type of manipulation will become useful later.
@@
(11) $\mathbf{v}^{\mathsf{H}}\mathbf{D}\mathbf{v}=-\mathbf{v}^{\mathsf{H}}\left({\frac{\mathbf{A}+\mathbf{A}^{\mathsf{T}}}{2}}\right)\mathbf{v}=-(\Re\lambda)(\mathbf{v}^{\mathsf{H}}\mathbf{v})\textbf{}$
($\mathsf{H}$ denoting Hermitian conjugate and $\Re$ denoting the real part), therefore $-\Re\lambda$ is bounded by the smallest and largest eigenvalues of $\mathbf{D}$ [14]. In addition, the trace of $\mathbf{D}\mathbf{C}^{-1}$ equals negative the trace of $\mathbf{A}$.
@@ 2.2. Comparison of experimental data with theoretical predictions
Now, we turn to the question of comparing experimental data and theoretical predictions under a linear model. As stated in the introduction, we may not necessarily care about statistically resolvable deviations if they are quantitatively small. We thus desire a way to evaluate the size of deviations. We may consider comparing the deviation between experiment and theory of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ to $\mathbf{C}$. That is, if we consider a whitened dataset for which $\mathbf{C}=\mathds{1}$, the entries of the “deviation matrix” (i.e., experiment minus theory of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$) in these coordinates should be compared to unity. In the following, we develop a coordinate-independent description of this procedure.
First, we consider the covariance matrix. We denote the deviation $\mathbf{M}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle_{\text{expt}}-\mathbf{C}_{\text{theo}}$, where “expt” and “theo” subscripts denote experimental and theoretical values, respectively. From now on we suppress the “theo” subscript and treat $\mathbf{C}$ as fixed. We may imagine an ensemble of stochastic systems, each having its own value of $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ which may differ from $\mathbf{C}$, and again denote the difference by $\mathbf{M}$. We assume that the ensemble mean of $\mathbf{M}$ is zero, and we now ask the question of what a suitable ensemble covariance of $\mathbf{M}$ would be that reflects significant deviations. We may then compare the elements of the experimental value of $\mathbf{M}$ to the square root of the ensemble variance. For biological applications, a factor of 0.3 or more may be considered significant.
@@
where $\sim$ denotes ensemble expectation. We may treat $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$, and multiply the left-hand side (l.h.s.) by the matrix inverse of the right-hand side (r.h.s.). We denote by $\bm{\Gamma}$ the resulting matrix:
@@
In coordinates where $\mathbf{C}=\mathds{1}$, the matrix $\bm{\Gamma}$ takes the form $\mathbf{v}\mathbf{v}^{\mathsf{T}}$, and therefore has rank at most 1. Thus, $\bm{\Gamma}$ has at most one non-zero eigenvalue, which is equal to its trace. We have:
(14) $\operatorname{tr}(\bm{\Gamma})=\Gamma^{ij}_{ij}=M^{ij}M^{i^{\prime}j^{\prime}}C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}=\operatorname{tr}((\mathbf{M}\mathbf{C}^{-1})^{2}).$
The second-rank mixed tensor $\mathbf{M}\mathbf{C}^{-1}$ has real eigenvalues because it is symmetric in “covariance-identity” coordinates. We see that the r.h.s. of the above is the sum of the squared eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$. We may use this quantity to evaluate quantitative significance [^5] It may be noticed that in “covariance-identity” coordinates, the “cross-covariances” ($M^{ij}$ where $i\neq j$) are counted with double the contribution to $\operatorname{tr}(\bm{\Gamma})$ as compared to “self-covariances” ($M^{ij}$ where $i=j$). This is necessary for a coordinate-independent evaluation. We may illustrate this on a two-dimensional example, with variables $x$ and $y$. We fix $\langle x^{2}\rangle=\langle y^{2}\rangle=1$, while $\langle xy\rangle$ is allowed to deviate from 0. Changing coordinates to $x^{\prime}:=(x+y)/\sqrt{2}$ and $y^{\prime}=(y-x)/\sqrt{2}$, we obtain $\langle{x^{\prime}}^{2}\rangle=1+\langle xy\rangle$, $\langle{y^{\prime}}^{2}\rangle=1-\langle xy\rangle$, and $\langle x^{\prime}y^{\prime}\rangle=0$. We thus see that if squared deviations of $\langle{x^{\prime}}^{2}\rangle$ and $\langle{y^{\prime}}^{2}\rangle$ to 1 are to be added, we must add the contribution due to $\langle xy\rangle$ twice. Furthermore, for a two-dimensional Ornstein–Uhlenbeck process with $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$, the time-averages for a trajectory of finite length, denoted by overlines, satisfy $\langle(\overline{x^{2}}-1)^{2}\rangle=2\langle(\overline{xy})^{2}\rangle$..
First, however, we must address a problem with Eq. (12), which is that it does not obey symmetry in the indices. We may remedy this by symmetrizing its r.h.s. To justify this choice, consider a complete set of left eigenvectors of $\bm{\Gamma}$. Such a set exists because a rank-one matrix can always be diagonalized. In dimension $d$, there are $d(d-1)/2$ linearly independent left eigenvectors such that $v_{ji}=-v_{ij}$; we call these “antisymmetric” left eigenvectors and are associated with a zero eigenvalue. We may assume that they are included our complete set of left eigenvectors. Let $v_{ij}^{(k)}$ be the remaining left eigenvectors of $\bm{\Gamma}$ with eigenvalues $\mu^{(k)}$ ($k$ taking $d(d+1)/2$ values), so that:
(15) $v_{ij}^{(k)}M^{ij}M^{i^{\prime}j^{\prime}}=\mu^{(k)}v_{ij}^{(k)}C^{ii^{\prime}}C^{jj^{\prime}}.$
@@
(16) $v_{ij}^{(k)}M^{ij}M^{i^{\prime}j^{\prime}}=\mu^{(k)}v_{ij}^{(k)}\frac{C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}}}{2}.$
Now if we neglect permutations in $(i,j)$ and $(i^{\prime},j^{\prime})$ [^6] I.e., only one of $(i,j)$ and $(j,i)$ is included in the list of “super-indices” (and similarly for $(i^{\prime},j^{\prime})$). This is necessary to invert $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$, since this quantity is invariant upon swapping $i\leftrightarrow j$ (or $i^{\prime}\leftrightarrow j^{\prime}$)., we can see that the matrix product of $M^{ij}M^{i^{\prime}j^{\prime}}$ with the matrix inverse of $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$, which we will denote $\bm{\Gamma}^{\prime}$, has left eigenvectors with elements $v_{ij}^{(k)}+v_{ji}^{(k)}$ for $i\neq j$ and $v_{ij}^{(k)}$ for $i=j$ with the same eigenvalues $\mu^{(k)}$. The resulting left eigenvectors of $\bm{\Gamma}^{\prime}$ are linearly independent, since if $\sum_{k}c^{(k)}(v_{ij}^{(k)}+v_{ji}^{(k)})=0$ for some scalars $c^{(k)}$, then $\sum_{k}c^{(k)}v_{ij}^{(k)}$ is antisymmetric and therefore $c^{(k)}=0$ by assumption of linear independence of the original set of left eigenvectors. Thus, aside from the zero eigenvalues of $\bm{\Gamma}$ associated with antisymmetric left eigenvectors, the eigenvalues of $\bm{\Gamma}$ and $\bm{\Gamma^{\prime}}$ are identical. We thus arrive at the relation:
@@
(18) $\operatorname{tr}(\bm{\Gamma})\sim\frac{d(d+1)}{2}.$
The r.h.s. is the number of degrees of freedom in $\mathbf{M}$, accounting for symmetry. To evaluate quantitative significance, we may compare the square roots of the above. We see that it is not suitable to compare the eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$ to unity, as there are only $d$ of these while the sum of their squares is compared to $d(d+1)/2$.
Now, we turn to covariance functions with a time-lag. We may again denote the deviation $\mathbf{M}:=\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{expt}}-\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{theo}}$. Now, there is no symmetry requirement for the components of $\mathbf{M}$, and hence we can take them as obeying Eq. (12). Collectively, if $\sqrt{M^{ij}M^{i^{\prime}j^{\prime}}C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}}$ is comparable to $d$, then $\mathbf{M}$ may be considered (collectively) quantitatively significant. In a sense, we are done. We may however consider what happens if we separate $\mathbf{M}$ into symmetric and antisymmetric components, denoted by $\mathbf{S}$ and $\mathbf{T}$ respectively:
@@
The r.h.s.’s are the sums of the squared moduli of the eigenvalues of $\mathbf{S}\mathbf{C}^{-1}$ and $\mathbf{T}\mathbf{C}^{-1}$. Also, straightforward calculation reveals that:
@@
The r.h.s. is $\operatorname{tr}(\bm{\Gamma})$, as before (although not $\operatorname{tr}((\mathbf{M}\mathbf{C}^{-1})^{2})$, since $\mathbf{M}$ is in general not symmetric). We have thus decomposed this into contributions from the symmetric and antisymmetric components of $\mathbf{M}$. Using Eq. (12), we then have:
@@
The ensemble variance of $\mathbf{S}$ is therefore the same as what was previously argued for $\mathbf{M}$ in the symmetric case. The ensemble variance of $\mathbf{T}$ is similar, except that it obeys antisymmetry rather than symmetry. The ensemble covariance between $\mathbf{S}$ and $\mathbf{T}$ vanishes, which makes sense because $\mathbf{S}$ and $\mathbf{T}$ acquire opposite signs upon time reversal. For “collective” quantitative significance, the above relations imply:
@@
The r.h.s.’s correspond to the number of degrees of freedom in $\mathbf{S}$ and $\mathbf{T}$, respectively. In this way, $\mathbf{S}$ and $\mathbf{T}$ may be individually evaluated for quantitative significance, if so desired. Similarly to before, it is not suitable to compare the eigenvalues of $\mathbf{S}\mathbf{C}^{-1}$ and $\mathbf{T}\mathbf{C}^{-1}$ to unity.
@@
2.3. Angular momentum and probability current density
The dynamics of the multivariate Ornstein–Uhlenbeck process is determined by the matrices $\mathbf{A}$ and $\mathbf{D}$. However, to alternatively characterize the system by quantities even and odd under time reversal, we may consider the angular momentum matrix:
@@ 2.3. Angular momentum and probability current density
2.3.1. Stochastic rotation frequencies
Now, we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. We will see that the possible values of $\omega_{\mathrm{stoch}}$ are restricted by the eigenvalues of $\mathbf{A}$. The angular momentum is $L:=\langle x\dot{y}-\dot{x}y\rangle$. It is related to the matrix $\bm{\Omega}$ by the relation:
@@ 2.3.1. Stochastic rotation frequencies
where we used the Lyapunov relation. Upon taking determinants, we get:
@@
(In the three-dimensional case, $\bm{\Omega}$ has at most a single pair of non-zero purely imaginary eigenvalues, and the stochastic rotation frequency is half the magnitude of the angular momentum vector evaluated in “covariance-identity” coordinates.)
The stochastic rotation frequency when $\mathbf{A}$ has two real eigenvalues has been computed in [15] and is easily derived from the above formula in coordinates where:
@@
2.3.2. Quantitative significance of broken detailed balance
At this point, we still do not have a good dimensionless measure of when broken detailed balance is quantitatively significant. To address this, we turn to a modified definition of gain matrix [13, 4]:
@@ 2.3.2. Quantitative significance of broken detailed balance
where the last equality follows from the Lyapunov equation. (We may understand the factor of 2 in the definition of $\mathbf{H}$ by considering that the symmetric part of $\mathbf{A}\mathbf{C}$ is $-\mathbf{D}$ by the Lyapunov equation, so the antisymmetric part should be compared to $\mathbf{D}$.) The entropy production rate $\dot{S}$ is related to $\mathbf{H}$ as $\dot{S}=-\operatorname{tr}(\mathbf{A}\mathbf{H})$ [13]. By a similar argument as in [13], in coordinates in which $\mathbf{D}=\mathds{1}$, we see that $\mathbf{H}$ is antisymmetric and therefore, like $\bm{\Omega}$, has purely imaginary complex conjugate eigenvalues. As an alternative derivation, following [3], we may consider that $\mathbf{D}^{-1/2}\mathbf{H}\mathbf{D}^{1/2}$ is antisymmetric and therefore $\mathbf{H}$ is similar to an antisymmetric matrix. In two dimensions the (dimensionless) eigenvalues ±$\pm ih$ are given by a simple formula similar to that of $\omega_{\mathrm{stoch}}$:
@@
If $h$ is at least comparable to unity (more precisely, $1/\sqrt{2}$; see end of subsubsection), then we consider that broken detailed balance is significant. We may illustrate this criterion in the cases of real or complex eigenvalues of $\mathbf{A}$. For two real eigenvalues $-\lambda_{x}$ and $-\lambda_{y}$ of $\mathbf{A}$ (Eq. (32)),
@@
For the case of complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ with $D^{xy}=0$ (Eq. (34)),
@@
We see that $h\geq\omega/\lambda$. Again, there are two possibilities for significant broken detailed balance. One possibility is that $\omega/\lambda$ is significant compared to unity. The other possibility is $\sqrt{D^{xx}}\ll\sqrt{D^{yy}}$ (or vice versa), which allows for the possibility that $\omega\ll\lambda$. Similarly to before, this corresponds to the case where $\mathbf{A}$ almost has a generalized eigenvector.
For angular momenta of $d$-dimensional systems, we may follow a similar line of reasoning as in section 2.2 but with symmetry replaced by antisymmetry, and with $\mathbf{C}$ replaced by $2\mathbf{D}$, which leads to the “ensemble covariance”:
@@
For a collective comparison, we have:
(44) $L^{ij}L^{i^{\prime}j^{\prime}}D^{-1}_{ii^{\prime}}D^{-1}_{jj^{\prime}}\sim 2d(d-1).$
Similarly to before, for high $d$ it is not suitable to compare the eigenvalues of $\mathbf{H}$ to unity. Furthermore, if the time-step is not small and angular momenta are measured using discrete time, then for the above comparisons, it may be suitable to use instead discrete-time estimators of the diffusion.